An investor placed her money in a venture paying 7% interest. If she had another $1500, she would have been able to invest her money at 10% and gain an additional $255 in annual interest. How much did she invest
Answer by mananth(16946)

An investor placed her money in a venture paying 7% interest. If she had another $1500, she would have been able to invest her money at 10% and gain an additional $255 in annual interest. How much did she invest
...
let her invest $x
x+1500 @ 10%
interest she would have gained = 255+0.07x
..
(x+1500)*0.1=225+0.07x
0.1x+150=225+0.07x
0.1x-0.07x=225-150
0.03x=75
x=75/0.03
x=2500 amount she invested
...
m.ananth@hotmail.ca